OpenStudy (bestpersonever123):

Joe opens a bank account, and starts with a deposit of $100. After that he deposits $20 each week. So what is the minimum number of weeks need for Joe in order to deposit $290?

OpenStudy (bestpersonever123):

is the answer 14 weeks?

OpenStudy (anonymous):

So we start with $100 + 20 = 120 + 20 = 140 + 20 = 160 + 20 = 180 + 20 = 200 + 20 = 220 + 20 = 240 + 20 = 260 + 20 = 280 + 20 = 300 (300 ≥ 290) (To find the answer simply add the number of weekly deposits) A quicker way to solve this is to take the final amount we want, subtract the initial amount (290 - 100 = 190) and divide by 20. Round your answer up to the nearest whole number and you have the answer.